a 3 kg cart starting from rest on a frictionless surface is pulled by a horizontal mass-less rope with a constant force of 6 N. What is the acceleration of the cart, and how fast is it moving at the end of two seconds?
1
Expert's answer
2012-10-19T12:04:46-0400
According to the second Newton's law, the acceleration of a cart is

a = F/m = 6[N]/3[kg] = 2[m/s²]

After two seconds the velocity of a cart is

V = at = 2[m/s²]*2[s] = 4[m/s].